36.) Two positive numbers differ by 11, and their square roots differ by 1. Find the numbers.
Solve by any method.
8.) x(x + 1) = 12


Find the complex solutions to the quadratic equations
12.) x^2+6x+10=0

Answer by RAY100(1637)

36) try 36, and 25
8) x(x+1)=12
x^2+x-12=0
(x+4)(x-3)=0
x= 3,-4
check
3(4)=12 ok
-4(-3)=12,,ok
12) x^2+6x+10=0
use quadratic formula
a=1,b=6,c=10
x= ( -6=/-sqrt(6^2-4(1)(10))/2(1)
x= (-6+/- sqrt(36-40))/2
x=-3=/-i
check
(-3+i)^2+6(-3+i)+10=0
9+i^2-6i+6i-18+10=0
but i^2=-1
8-18+10=0,,,,ok